OpenStudy (anonymous):

Need help understanding a small part of this integration problem. Give me a minute to type it.

OpenStudy (anonymous):

\[2\int\limits_{0}^{4}\frac{ 3x }{ 1+x^2 }\]

OpenStudy (anonymous):

so if i let u - 1+x^2.. then du = 2x dx. How do I account for the other x

OpenStudy (dan815):